OpenStudy (anonymous):

Evaluate. (x+3)/(x-6)-(x^2+6x)/(x^2-36)

OpenStudy (lalaly):

it's easy just give it a try its all about factorisation

OpenStudy (anonymous):

i factored..but i keep coming up with 3/x-6

OpenStudy (lalaly):

\[\large{\frac{(x+3)}{(x-6)}-\frac{x^2+6x}{x^2-36}}\]\[\large{\frac{x+3}{x-6}-\frac{x \cancel{(x+6)}}{(x-6) \cancel{(x+6)}}}\]\[\large{\frac{x+3-x}{x-6}=\frac{3}{x-6}}\] YOU're ANSWE IS RIGHT:D

OpenStudy (anonymous):

thanks